OpenStudy (anonymous):

Given that h(x)=x^2g(x) and g(3)=5 and g'(3)=7 find h'(3)

OpenStudy (anonymous):

You can use the product rule. h'(x) = f(x)g'(x) + g(x)f'(x) where in this problem f(x) = x²

OpenStudy (anonymous):

so f(x)=x^2 and f'(x)=2x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

then evaluate them both at 3

OpenStudy (anonymous):

what about g(x) and g'(x)?

OpenStudy (anonymous):

well they're asking for h'(3), so you don't need g(x) and g'(x) since they gave you the values at 3.

OpenStudy (anonymous):

so how would you use the product rule?

OpenStudy (anonymous):

Above is the product rule generally. Specifically at x = 3. \[h'(3) = f(3)g'(3) + g(3)f'(3)\]

OpenStudy (anonymous):

g(3) and g'(3) are given. f(3) and f'(3) are found by evaluating x² and 2x at 3 respectively

OpenStudy (anonymous):

oh is it h'(3)=9*7+5*6

OpenStudy (anonymous):

f(3) = 3² = 9 f'(3) = 2(3) = 6

OpenStudy (anonymous):

yep exactly

OpenStudy (anonymous):

so the answer is 93

OpenStudy (anonymous):

thanks for your help peachpi

OpenStudy (anonymous):

yes that's right. You're welcome